OpenStudy (anonymous):

Use the function rule to complete the table. -10x+y=4 X: -2, -1, 0 ,1, 2 Y: __, __, __, __,__

OpenStudy (anonymous):

Solve for y in -10x+y = 4.

OpenStudy (anonymous):

I dont know how to do it, im confused

OpenStudy (anonymous):

i'll try my best

OpenStudy (anonymous):

\[-10x+y=4\] if \(x=-2\) you get \[-10\times -2+y=4\\ 20+y=4\\ y=-16\]

OpenStudy (anonymous):

see you are getting help

OpenStudy (anonymous):

i am?

OpenStudy (anonymous):

if \(x=-1\) you have \[-10\times -1+y=4\\10+y=4\\ y=-6\]

OpenStudy (anonymous):

etc

OpenStudy (jdoe0001):

\(\large -10{\color{red}{ x}}+y=4\qquad \qquad \begin{array}{ccllll} x\qquad &y\\ \hline\\ -2&-10({\color{red}{ -2}})+y=4\\ -1&-10({\color{red}{ -1}})+y=4\\ 0&-10({\color{red}{ 0}})+y=4\\ 1&-10({\color{red}{ 1}})+y=4\\ 2&-10({\color{red}{ 2}})+y=4 \end{array}\)